OpenStudy (iwanttogotostanford):

Solve 4^(2x) = 7^(x−1).

OpenStudy (iwanttogotostanford):

@phi

OpenStudy (phi):

are you studying logarithms ?

OpenStudy (iwanttogotostanford):

@phi yes

OpenStudy (phi):

have you learned how to "take the log" of both sides ?

OpenStudy (iwanttogotostanford):

no @phi

OpenStudy (phi):

"take the log" means write log in front of a number (or expression) for example, to take the log of 10 you write log 10

OpenStudy (iwanttogotostanford):

oh ok so, 4 would be log 4

OpenStudy (phi):

so log of 4 is log 4 write down your equation after you "take the log" of both sides (don't simplify anything, just write log on each side)

OpenStudy (iwanttogotostanford):

ok so I have: log 4^(2x)= log 7 ^(x-1)

OpenStudy (phi):

yes, and technically we should put in parens because we are doing \[ \log\left( 4^{2x}\right) = \log\left(7^{(x-1)} \right)\]

OpenStudy (iwanttogotostanford):

oh ok, got it! Now what?

OpenStudy (phi):

there is a "property" of logarithms that we now use (this is the reason we are using logs) the property is \[ \log(a^b) = b \log(a) \]

OpenStudy (iwanttogotostanford):

ok!

OpenStudy (phi):

any idea how to use that property on both sides of your equation? for example, if you apply it to the left side, what do you get ?

OpenStudy (iwanttogotostanford):

on the left side I will have: Log (4^2)=^2log(4) ?

OpenStudy (phi):

you mean \[ \log(4^2) = 2 \log(4) \] but you don't have 4^2 you have 4^(2x)

OpenStudy (phi):

try again

OpenStudy (iwanttogotostanford):

yes

OpenStudy (iwanttogotostanford):

so log(4^2) = 2log(4)

OpenStudy (iwanttogotostanford):

i put it in my calculator and I got 1

OpenStudy (iwanttogotostanford):

so would 1 be the answer

OpenStudy (iwanttogotostanford):

these are my answer choices: 2.35389 −2.35389 −1 1

OpenStudy (phi):

no. we'll get back to using the calculator later. but the left side is \[ \log(4^{2x} ) \]

OpenStudy (iwanttogotostanford):

ok!

OpenStudy (phi):

use the rule on that

OpenStudy (iwanttogotostanford):

how would I use the rule for log (4^2x)? Would I do: log (4*2)=2log(4) log(ab)=blog(a)

OpenStudy (iwanttogotostanford):

a= 4 and b=2

OpenStudy (phi):

the idea is to "match up" your problem with the rule